dukedadon:

how to solve this linear algebraic equation 2.4x + 1.4 = 3.8

06natrod52378:

Step one)2.4x=-1.4+3.8 Step two)-1.4+3.8 Step three)2.4x=2.4 Step four) x=1 i think im 99.9% sue of it if this is what i think it is

littleraw27:

smart

dukedadon:

bro, its sad to see that it's 6 of you guys viewing and only one person answered, why are yall just staring at the question???

Astrid1:

\[2.4x + 1.4 = 3.8\] Multiply both sides by 10: \[2.4x \times\:10+1.4 \times\:10=3.8\ \times\:10\] Refine: \[24x+14=38\] Subtract 14 by each side: \[24x+14-14=38-14\] Simplify: \[24x=24\] Divide each side by 24: \[\frac{24x}{24}=\frac{24}{24}\] Simplify: \[x=1\] So your answer should x=1.

06natrod52378:

@dukedadon wrote:
bro, its sad to see that it's 6 of you guys viewing and only one person answered, why are yall just staring at the question???
bc im the olny one with the balls to be wrong if i am

Astrid1:

I was literally typing all of that out god.

06natrod52378:

i said the same thing astrid u just copyed me

Astrid1:

I literally know how to do it, I just explained it further.

Astrid1:

Just like how I do with every math post I answer.

06natrod52378:

smhh

littleraw27:

stop arguing

06natrod52378:

wait i hate u i shouldnt be talking with u

dukedadon:

lmaoooo guys guys, relax.

littleraw27:

its all over your question